Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: Move to A configuration began this morning; 4-band dipoles in place through the 15th (ongoing issues as described last week); correlator good to 63/64 BlB pairs. Some ongoing issues with missing first scans (recoverable). Report from AAS. Summer students begin arriving. PEPs due for science staff.
  • Key dates:
    • 07 May - 15 Jun: 4-band science observing
    • 10 Jun - 12 Sep: A configuration (move from 31 May - 09 Jun).
  • Staff Travel:
    • LS: 9 May - 20 Jun
    • GM: 25 May - 24 Jun
    • FO: 26 May - 09 Jun
    • FO: 12 Aug - 12 Sep
  • Visitors:
    • Robert Harris (AC982): 24 Apr - 16 Jul
    • Bob Sault: 16 May - 23 Jun

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Antennas: EA18 pushed to 6 Jun (rolled into A configuration move). EA21 next in.
    • Receivers: EA05 Ku; EA12 X due this week
    • Calibration (Perley): Ongoing investigation into the switched power compression/expansion.
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): 3-bit testing review (Dhawan). V stokes data taken (high frequency data pushed). Bent-pin; BlB shuffle ready for this afternoon (2 hours for Kerry).
  • Computing (Robnett):

Software status and updates (Butler)

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Configuration move week (work around these activities:
  • Mon
    • Holiday
  • Tue: Maintenance day:
    • 0630-? Testing/software development time
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Fri (not-yet-a-holiday but Monday is):
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags).
  • Correlator: [Clark/Carlson] Delay rocking across bands. Implementation pending factor-of-2 reconciliation.
  • Correlator: [Rupen/Carlson w/ Momjian/Sault] V stokes issues.
  • Array: [Gustaaf] Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).
  • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).

Testing This week

  • CM/CBE update testing (subarrays+ deconfiguration problems).
  • 4-band testing (Owen): further tests of modified subband/requantizer gains (Clark)
  • phased array (Sowinski/McMullin): further tests from Amy/Ken (as above)
  • Array: *; Individual issue sleuthing (focus on phase jumps).
  • High Frequency (Perley): Fast switching duty cycle tests.
  • Astrometry (Bruthaler): UVW precision tests.
  • Correlator: Rupen: S-band shadow testing
  • Correlator: Durand/McMullin/Pokorny/Butler; review of CRC error events (may involve testing).
  • Spectral line: Momjian: HI High resolution bandpass testing
  • 3-bit samplers (Dhawan+): lab testing this week.

  • V stokes testing (Sault/Momjian) TSPE0008_sb1850345)Complete; high frequency test pushed.
  • Spectral Line (Linz/Ott): Spectral dynamic range test (TSPE0010.3923557). Complete.
  • Spectral Line (Goss/Chomiuk): BP stability/repeatability (TSPE0005.3699117). Complete..

Commissioning Groups

  • Big deadline is end of this month for incorporation into links for Call for Proposals due 1 July.

    • Adv: Low band (Owen):
      • Notes:

RSRO/ECSO Reports

RSRO B configuration status: https://safe.nrao.edu/wiki/bin/view/EVLA/2011BRSROStatus

Individual reports (updates)

Misc:

  • RSRO Calendar: https://safe.nrao.edu/wiki/bin/view/EVLA/RSROCalendar
  • RSRO DnC, C config status: https://safe.nrao.edu/wiki/bin/view/EVLA/2010CRSROStatus
  • Test Project Codes == *TRSR*
  • Guidelines for RSRO observing: https://safe.nrao.edu/wiki/bin/view/EVLA/RSROObservingPreparationGuidelines
  • Notes:
    • Current Residents
      • D.J. Saikia (10C-119/AI145)::RSRO-75; 300G(2); x7106-POC: Momjian (Low Frequency)
      • J. Irwin (10C-119AI145)::RSRO-75; 300G(1); x7333 -- POC: Myers (Polarimetry)
      • Peter Hofner (10B-124/AH1011)::RSRO-26; 200G(1); x7496 -- POC: Ott (CASA testing)
      • Sergio Dzib (10B-131); 200G;x7496 -- POC: Mioduszewski (Astrometry)
      • Jacqueline van Gorkom (10C-211/AV324); 200D/x7098 -- POC: Momjian (Low Frequency)
    • Upcoming
      • 06 Jun - 31 May (2012): Andreas Brunthaler::200E/x7246 -- POC: Mioduszewski (Astrometry/VLBI)
      • 12 Jun-10 Aug: Jane Greaves::(AC982)::RSRO-4; 300G;x7106 -- POC: Chandler (High Frequency)
      • 12 Jun - 25 Jun: Laura Chomiuk (*); 200B/x7497 -- POC: Goss (Spectral Line)

Demo Science Status

AOB

Misc Actions

  • Old/Low priority
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)

-- JosephMcMullin - 2011-05-31
Topic revision: r2 - 2011-05-31, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback